OpenStudy (toxicsugar22):

According to the following reaction, how many grams of water are necessary to form 0.478 moles oxygen gas? 2 H2O (l) = 2 H2 (g) + O2 (g)

OpenStudy (chmvijay):

for 1 mole of O2 2*18 =36 g of H2O is required for 0.478 moles of O2 =36*0.478/1 =17.20 mole of O2

OpenStudy (chmvijay):

where is reaction dudee

OpenStudy (chmvijay):

hey dude its not the equation i think

OpenStudy (chmvijay):

28 g of N2 requires = 6 g of H2 29.9 gram of N2 =6*29.9/28 =6.40

OpenStudy (chmvijay):

u try it i have solved already the two

OpenStudy (anonymous):

normally such questions require to look for the limiting reagent in a case when none of the moles have been given in this case since the moles of oxygen have been given then we look for the ratio of oxygen to water in the reaction now we have a 1:2 molecular ratio of oxygen gas to water in the reaction as you can see.that is you see that 1 oxygen in the reaction give rise to 2 water that is 2H20=O2 as you wrote this means if 1O2=2H2O then 0.478=XH2O which is 2*0.478 moles now with this you simply know that moles =mass /molar mass now you have the moles of 2 molecules of water so now you need the molar mass of these H2O molecules which is simply 18 because the 2 molecules of water have been used to get the ratio so no need to multiply 18 by 2 anymore now that you have moles for water and molar mass of the water in the reaction you put them in the formula above which means 18*2*0.478=Ans

OpenStudy (toxicsugar22):

According to the following reaction, how many grams of water will be formed upon the complete reaction of 21.6 grams of oxygen gas with excess hydrogen sulfide? 2 H2S (g) + 3 O2 (g) = 2 H2O (l) + 2 SO2 (g)

OpenStudy (toxicsugar22):

can you help me solve this

OpenStudy (anonymous):

excess hydrogen sulfide means that oxygen is the limiting reagent so you always need to know this first now even when you know this the calculations help you by giving you the mass of some molecules such that you do not need to stress yourself about which molecule or reagent to use like this one but mind you if you have the mole of both reagents you need to know the limiting reagent i ll try to look for a way to let you understand that better but do not worry about it for now focus on what i am about to do only we need to find the moles of oxygen gas alone first from my approach and that is given by Mass/Molar Mass and that is going to be 21.6/32=Ans now after this you use my molecular ratio approach which is if 3O2=2H2O then 21.6/32=XH20 when you get this mole of H20 you are really fine to go ahead you just come back to your simple formula moles =mass/molar mass which means mass of H2O=moles above*18 grams just take some time to view the steps see this you can either

OpenStudy (toxicsugar22):

thanks

OpenStudy (anonymous):

you can either go the fast way which is harder to understand without some good algebra skills or you go this way

OpenStudy (toxicsugar22):

thansk I got it know